# Trajectory Consistency Distillation

### v1

@@ 4.1 Trajectory Consistency Function
The distillation error from the consistency model can be reduced by expanding the original boundary conditions to encompass the entire trajectory. To facilitate this, we introduce the Trajectory Consistency Function (TCF), designed to enable comprehensive tracking along the full trajectory.
${\bm{f}}({\bm{x}}_{t},t,s)\mapsto{\bm{x}}_{s}.$ (13)
The trajectory consistency function possesses the property of trajectory consistency, enhancing the original self-consistency property in an endpoint-unrestricted trajectory. Specifically, its output remains consistent for arbitrary sets $({\bm{x}}_{t},t)$ with the given $s$ that belong to the same PF ODE trajectory, where $0\leqslant s\leqslant t\leqslant T$,
${\bm{f}}({\bm{x}}_{t},t,s)={\bm{f}}({\bm{x}}^{\prime}_{t},t^{\prime},s)\quad\forall t,t^{\prime},s\in[0,T].$ (14)
Figure 3: Synthesis results across various NFEs. Due to accumulated errors in multistep sampling, LCM experiences a loss of image detail, leading to a degradation in performance, whereas TCD addresses this issue. Additional samples are available in Section E.1.
@@
The semi-linear structure of the empirical PF-ODE revealed by (Lu et al., 2022a, b; Zhang & Chen, 2022) motivates us to parameterise the trajectory consistency function using the exponential integrators form, as shown in Equation 15.
${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\sigma_{s}\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\hat{{\bm{x}}}_{\bm{\theta}}(\hat{{\bm{x}}}_{\lambda},\lambda)\text{d}\lambda,$ (15)
where $\lambda_{t}:=\log(\alpha_{t}/\sigma_{t})$ is the log-SNR, and ${\bm{x}}_{\bm{\theta}}$ is a trainable network with parameter $\bm{\theta}$.
@@
$\begin{split}{\bm{f}}_{\bm{\theta}}&({\bm{x}}_{t},t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\\ &\sigma_{s}\sum_{n=0}^{k-1}{\bm{x}}_{\bm{\theta}}^{(n)}(\hat{{\bm{x}}}_{\lambda_{t}},\lambda_{t})\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\frac{(\lambda-\lambda_{t})^{n}}{n!}\text{d}\lambda+\mathcal{O}(h^{k+1}),\end{split}$ (16)
@@
Here, we consider 1st-order and 2nd-order estimations, omitting high-order terms $\mathcal{O}(h^{k+1})$.
@@
$\begin{split}\text{TCF(2):}\quad{\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},&t,s)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}-\alpha_{s}(e^{-h}-1)\\ &\left((1-\frac{1}{2r})\hat{\bm{x}}_{\theta}({\bm{x}}_{t},t)+\frac{1}{2r}\hat{\bm{x}}_{\theta}(\hat{\bm{x}}_{u},u)\right),\end{split}$ (18)
@@
Broadening the Boundary Condition.
The boundary condition in the CM is limited to intervals extending from any start point on the solution trajectory to the origin. Conversely, TCF mitigates this constraint, allowing the model to handle any interval along the PF-ODE trajectory and preventing trivial solutions ${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},t,s)\equiv 0$ from arising in TCD training. Consequently, we broaden the boundary condition to encompass a more comprehensive range of trajectory intervals,
${\bm{f}}_{\bm{\theta}}({\bm{x}}_{s},s,s)={\bm{x}}_{s},$ (20)
it is evident that all three types of our parametrisation satisfy the broad boundary condition effortlessly.
@@
Thus, we could express the object of trajectory distillation in alignment with reconstruction:
@@
Below, we offer a theoretical justification based on asymptotic analysis to explain how trajectory consistency distillation optimizes the distillation error.
Theorem 4.1.
Let $\Delta t:=\rm{max}_{n\in[\![1,N-1]\!]}\{\lvert t_{n+1}-t_{n}\rvert\}$, and $\bm{f}(\cdot,\cdot,\cdot;\bm{\phi})$ be the trajectory consistency function of the empirical PF ODE in Equation 3. Assume $\bm{f}_{\bm{\theta}}$ satisfies the Lipschitz condition, that is, there exists $L>0$ such that for all $t\in[0,T]$, $\bm{x}$ and $\bm{y}$, we have $\left\lVert{\bm{f}_{\bm{\theta}}(\bm{x},t,s)-\bm{f}_{\bm{\theta}}(\bm{y},t,s)}\right\rVert_{2}\leqslant L\left\lVert{\bm{x}-\bm{y}}\right\rVert_{2}$. Assume further that for all $n\in[\![1,N-1]\!]$, the $p$-th order ODE solver called at $t_{n+1}$ has local error uniformly bounded by $\mathcal{O}((t_{n+1}-t_{n})^{p+1})$ with $p\geqslant 1$. Then, if there is a $\bm{\theta^{*}}$ so that $\mathcal{L}_{\rm{TCD}}^{N}(\bm{\theta^{*}},\bm{\theta^{*}};\bm{\phi})=0$, for any $n\in[\![1,N-1]\!]$ and $m\in[\![1,n]\!]$, we have
$\begin{split}\mathop{\textnormal{sup}}\limits_{n,m,{\bm{x}}}\|{\bm{f}}_{\bm{\theta^{*}}}({\bm{x}},t_{n},t_{m}),{\bm{f}}({\bm{x}},&t_{n},t_{m};\phi)\|_{2}\\ &=\mathcal{O}\left((\Delta t)^{p}\right)(t_{n}-t_{m}).\end{split}$
Proof.
The proof is provided in  Section D.2. ∎
Theorem 4.1 implies that the distillation error of TCF is upper bound by that of CD presented in Equation 7.
@@ 4.2 Strategic Stochastic Sampling
The proposed trajectory consistency function not only optimizes the distillation loss but also enables the model to access non-origin destinations along the PF ODE. This capability allows Strategic Stochastic Sampling (SSS) to further reduce discretisation errors and estimation errors introduced in each sampling step.
@@
Specifically, every sampling step in SSS includes the denoise sub-step according to the ODE solver and the diffuse sub-step based on Langevin SDE. In comparison with multistep consistency sampling, where the endpoint and noise level are fixed, SSS introduces the additional parameter $\gamma$ to control the destination point for the denoise step and allows for the adjustment of random noise level for the diffuse step, as detailed in Figure 1(b) and Algorithm 4. This parameter is referred to as the stochastic parameter.
In the denoising sub-step, our focus is on reducing the errors that arise from the $n$-th step in Equation 11 to $\mathcal{O}(\sqrt{\tau_{n}-(1-\gamma)\tau_{(n+1)}})$ when $n\in[\![1,N-1]\!]$, by the prediction of the non-origin, and ultimately optimize the accumulated error as shown in Theorem 4.2.
Theorem 4.2.
As the strategic stochastic sampling process defined in Algorithm 4, and denoting the $N>1$ steps sampling results with trained trajectory consistency model ${\bm{f}}_{\bm{\theta^{*}}}$ as $q_{\bm{\theta^{*}},N}={\bm{f}}_{{\bm{\theta^{*}}},\tau_{N}}\sharp\mathcal{N}(\alpha_{\tau_{N}}{\bm{x}}_{\tau_{(N-1)}\rightarrow 0},\sigma_{\tau_{N}}\bm{I})$, the results with exact trajectory consistency function ${\bm{f}}_{\bm{\theta}}$ as $p_{N}$, then the accumulated error in SSS is defined as:
$TV(q_{\bm{\theta^{*}},N},p_{N})=\mathcal{O}\left(\sum_{n=1}^{N-1}\sqrt{\tau_{n}-(1-\gamma)\tau_{n+1}}+\sqrt{\tau_{N}}\right),$ (23)
where $\gamma\in[0,1]$ is the parameter controlling the destination in each denoising step.
Proof.
The proof is provided in Section D.3. ∎
While sufficient stochasticity in diffuse sub-step helps reduce both the estimation errors and discretisation accumulated in earlier sampling steps and drives the sample towards the desired marginal distribution, a similar finding is also observed in (Karras et al., 2022; Xu et al., 2023b). In SSS, we encourage the step size of the diffuse sub-step to be smaller than that of the denoise sub-step, which is contrary to (Xu et al., 2023b). It is also worth noting that when $\gamma$ is low, the estimation error plays a more important role. Hence, the optimal value of $\gamma$ should be determined empirically, as we show in Section 5.3.
@@
Figure 4: Qualitative comparison. For each prompt, images are generated using the same random seed for every model, without any cherry-picking. More results are provided in Section E.3.
@@
Figure 5: Qualitative effects of stochastic parameter $\gamma$ with same NFEs. Images under the same prompt but with different $\gamma$ applied during sampling. The leftmost image is sampled from LCM (Luo et al., 2023b). More samples can be found in Section E.2.
@@
4.3 Extension to Large Text Conditional Models
Trajectory consistency distillation can directly occur on top of the parameters of a pre-trained diffusion model as the fine-tuning process. For scaling TCD to larger models (e.g., SDXL) with significantly reduced memory consumption, we incorporate Low-Rank Adaptation (LoRA) (Hu et al., 2021), a parameter-efficient fine-tuning method, into the distillation process. Additionally, the parameters of LoRA can be identified as a versatile acceleration module applicable to different fine-tuned models or LoRAs sharing the same base model without the need for additional training, aligning with the observations in (Luo et al., 2023b).
5 Experiments
5.1 Experimental Setup

### v2

@@ 4 Trajectory Consistency Distillation
4.1 Background
Given the $\epsilon$-prediction model $\epsilon_{\bm{\theta}}(\cdot)$, Luo et al. (2023a) used DDIM (Song et al., 2020a) to re-parameterize the consistency function in (Song et al., 2023) but still followed the skip connection form as in Equation 11 with boundary conditions $c_{\text{skip}}(0)=1$ and $c_{\text{out}}(0)=0$. However, such a parameterization process includes a discretization error caused by the finite difference approximation of an explicit numerical method (Song et al., 2020a) with the spacing $t\rightarrow 0$. Besides, in practical experiments with LCM released training codes  [^1] https://github.com/luosiallen/latent-consistency-model, we noticed the implementation of preconditions $c_{\text{skip}}$ and $c_{\text{out}}$ only matches the boundary condition but stray from the initial definition in (Karras et al., 2022).
In order to better fit the formulation in LDM (Rombach et al., 2022) and SDXL (Podell et al., 2023), we first dispose of the skip-connection form with preconditions $c_{\text{skip}}$ $c_{\text{out}}$ and rewrite Equation 11 as:
$\begin{split}\tilde{{\bm{f}}}_{\bm{\theta}}({\bm{x}}_{t},t)&=\frac{{\bm{x}}_{t}-\sigma_{t}\hat{\epsilon}_{\bm{\theta}}({\bm{x}}_{t},t)}{\alpha_{t}}\\ &=\frac{1}{\alpha_{t}}{\bm{x}}_{t}-\frac{\sigma_{t}}{\alpha_{t}}\hat{\epsilon}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t)\\ &=\frac{\alpha_{0}}{\alpha_{t}}{\bm{x}}_{t}-\alpha_{0}\left(\frac{\sigma_{t}}{\alpha_{t}}-\frac{\sigma_{0}}{\alpha_{0}}\right)\hat{\epsilon}_{\bm{\theta}}(\hat{{\bm{x}}}_{t},t),\end{split}$ (14)
which is in the form of DDIM (Song et al., 2020a). As we mentioned before, this function also includes an approximation error with huge spacing.
Lu et al. (2022a) pointed out Equation 14 is actually an approximated case for the exact solution of PF-ODEs with semi-linear structure with an exponential integrator in terms of $\bm{\epsilon}_{\bm{\theta}}$:
$\hat{\bm{x}}_{s}=\frac{\alpha_{s}}{\alpha_{t}}{\bm{x}}_{t}+\alpha_{s}\int_{\lambda_{t}}^{\lambda_{s}}e^{-\lambda}\hat{\bm{\epsilon}}_{\bm{\theta}}(\hat{{\bm{x}}}_{\lambda},\lambda)\text{d}\lambda,$ (15)
and analyzed the corresponding error associated with $\lambda_{t}$ and $\lambda_{s}$, characterized as $\mathcal{O}((\lambda_{s}-\lambda_{t})^{k})$ with $k$-th Taylor expansion. In this context, $\bm{\epsilon}_{\bm{\theta}}$ is a trainable network with parameter $\bm{\theta}$, and $\lambda_{t}:=\log(\alpha_{t}/\sigma_{t})$ is the log-SNR. We note that when using the same order $k$ of Taylor expansion, the error decreases as the interval between $s$ and $t$ reduces. Another point worth noting is that shortening the time intervals makes the boundedness assumption in (Lu et al., 2022a) easier to meet, which ensures the error correspondingly decreases as the order $k$ increases. We would like to highlight that such an mentioned error is the discretization error arising during the parameterisation process in  Equation 11. To further improve the guided sampling quality, another exact solution of PF-ODEs with ${\bm{x}}_{\bm{\theta}}$ is proposed in (Lu et al., 2022b):
$\hat{{\bm{x}}}_{s}=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\sigma_{s}\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\hat{{\bm{x}}}_{\bm{\theta}}(\hat{{\bm{x}}}_{\lambda},\lambda)\text{d}\lambda,$ (16)
which leads to better guided sampling performance and shares analogous properties with Equation 15. However, given that the consistency function is devoid of any guidance sampling, these two equations are equally valid views and can be used interchangeably (Kingma et al., 2021). In this work, we opt for the ${\bm{x}}_{\bm{\theta}}$ representation due to its greater intuitiveness, despite the fact that $\bm{\epsilon}_{\bm{\theta}}$ is mostly used for parameterization in practice.
Furthermore, Equation 9 also reveals that in the distillation process, the upper bound of the CD loss is not only related to the number of steps $\Delta t$ derived by solvers, but is also associated with the time interval $t\rightarrow 0$ spanned in the consistency function.
Figure 3: Synthesis results across various NFEs. Due to accumulated errors in multistep sampling, LCM experiences a loss of image detail, leading to a degradation in performance, whereas TCD addresses this issue. Additional samples are available in Section D.1.
@@ 4.2 Trajectory Consistency Function
Such errors as mentioned lead us to seek a better form for the consistency function by broadening the endpoint from the origin to a midpoint $s$ between 00 and $t$ within the semi-linear structure.
Recall Equation 6 and the definition of the consistency function in (Song et al., 2023): ${\bm{f}}({\bm{x}}_{t},t)\mapsto{\bm{x}}$, and the definition of the self-consistency property of the consistency function  Equation 5: The outputs of the consistency function are consistent for arbitrary pairs that belong to the same PF-ODE trajectory. The goal of consistency distillation is to estimate this consistency function and ensure the self-consistency property.
@@
We introduce the Trajectory Consistency Function (TCF) to re-parameter the output of ${\bm{x}}_{\bm{\theta}}(\cdot)$ to ${\bm{x}}_{s}$:
${\bm{f}}^{\rightarrow s}({\bm{x}}_{t},t)\mapsto{\bm{x}}_{s}.$ (17)
Regarding the selection of $s$, our objective is to ensure that $s$ conforms to the minimum error across various discrete situations; thus, $s$ is designated as an arbitrary integer within the range $[0,t]$.
And TCF also supports self-consistency. TCF remains consistent for arbitrary sets $({\bm{x}}_{t},t)$ with the given $s$ that belong to the same PF ODE trajectory:
${\bm{f}}^{\rightarrow s}({\bm{x}}_{t},t)={\bm{f}}^{\rightarrow s}({\bm{x}}_{t^{\prime}},t^{\prime})\quad\forall t,t^{\prime},s\in[0,T].$ (18)
@@
The semi-linear structure with the exponential integrators form (Lu et al., 2022a, b; Zhang & Chen, 2022) can be easily used to parameterise the trajectory consistency function,
[…]
@@
$\begin{split}{\bm{f}}^{\rightarrow s}_{\bm{\theta}}&({\bm{x}}_{t},t)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}+\\ &\sigma_{s}\sum_{n=0}^{k-1}\hat{\bm{x}}_{\bm{\theta}}^{(n)}(\hat{{\bm{x}}}_{\lambda_{t}},\lambda_{t})\int_{\lambda_{t}}^{\lambda_{s}}e^{\lambda}\frac{(\lambda-\lambda_{t})^{n}}{n!}\text{d}\lambda+\mathcal{O}(h^{k+1}),\end{split}$ (20)
@@
Here, we consider 1st-order and 2nd-order estimations, omitting high-order terms $\mathcal{O}(h^{k+1})$ in equation expression.
@@
$\begin{split}\text{TCF(2):}\quad\tilde{{\bm{f}}}^{\rightarrow s}_{\bm{\theta}}({\bm{x}}_{t},&t)=\frac{\sigma_{s}}{\sigma_{t}}{\bm{x}}_{t}-\alpha_{s}(e^{-h}-1)\\ &\left((1-\frac{1}{2r})\hat{\bm{x}}_{\theta}({\bm{x}}_{t},t)+\frac{1}{2r}\hat{\bm{x}}_{\theta}(\hat{\bm{x}}_{u},u)\right),\end{split}$ (22)
@@
According to Theorem 3.2 in (Lu et al., 2022a), the upper error bounds for TCF parameterisation is the same as DPM-Solver-$k$. Let ${\bm{f}}_{\bm{\theta}}^{\rightarrow s}({\bm{x}}_{t},t)$ be the ideal trajectory consistency function with exponential integrators form of the empirical PF-ODE, where $t$ and $s$ are the start timestep and end timestep along PF-ODE, we have the discretisation error regarding to the finite $k$-th order of TCF:
$\|\tilde{{\bm{f}}}_{\bm{\theta}}^{\rightarrow s}({\bm{x}}_{t},t)-{\bm{f}}_{\bm{\theta}}^{\rightarrow s}({\bm{x}}_{t},t)\|_{2}=\mathcal{O}(h_{\Delta t}^{k+1}),$ (23)
where $h:=\lambda_{s}-\lambda_{t}$ and $k$ is the TCF order. As identified in  (Zhang & Chen, 2022; Lu et al., 2022a), employing semi-linearity in diffusion ODEs results in minimal error.
@@
In order to unify the representation of TCF, we abuse
${\bm{f}}({\bm{x}}_{t},t,s):={\bm{f}}^{\rightarrow s}({\bm{x}}_{t},t),$ (25)
where $s$ can be a parameter either in the reparameterised process or the network.
We observe that TCF(S+) integrates the parameter $s$ into the network, yielding a formulation akin to the $G$ function in  (Kim et al., 2023), albeit with a different coefficient. The underlying reason is that both instances employ a neural network to directly get the integral term from $t$ to $s$. The key distinction lies in the dependency of CTM’s $G_{\theta}$ on the Euler Solver and the requirement of a ${\bm{g}}_{\theta}$ for estimating $\mathbb{E}[{\bm{x}}|{\bm{x}}_{t}]$. Conversely, TCFs adopt a semi-linear structure for a better consistency function and use the finite difference approximation to parameterize the integral term.
Broaden Boundary Conditions
The boundary condition in the CM is limited to intervals extending from any start point on the solution trajectory to the origin. And TCF is required to adhere to classical boundary conditions (Song et al., 2020b)
${\bm{f}}({\bm{x}}_{0},0,0)={\bm{x}}_{0},$ (26)
but also needs to meet a broaden boundary conditions
${\bm{f}}({\bm{x}}_{s},s,s)={\bm{x}}_{s},$ (27)
to encompass a more comprehensive range of trajectory intervals with arbitrary sets.
It is evident that our parametrisation satisfy both the classical boundary conditions and the broaden boundary condition effortlessly.
@@
Thus, we could express the object of trajectory distillation in alignment with reconstruction, similar to the consistency distillation loss in (Song et al., 2023):
@@
This objective function inherit the form in (Luo et al., 2023a; Song et al., 2023), we expand it to a broaden condition according to the definition of TCF in Equation 17 and ${\bm{f}}_{\bm{\theta}}({\bm{x}}_{t},t,s)$ still holds the form of ${\bm{f}}^{\rightarrow s}_{\bm{\theta}}({\bm{x}}_{t},t)$, which can be easily reparameterised and lead to a fast convergence. Such parameterisation of consistency function in semi-linear is pivotal for fast convergence and better results.
From Theorem 1 in (Song et al., 2023), we could found trajectory consistency distillation with a shorter spacing can reduce the distillation error from $\mathcal{O}\left((\Delta t)^{p}\right)(t_{n}-t_{0})$ to $\mathcal{O}\left((\Delta t)^{p}\right)(t_{n}-t_{m})$.
It is observed that the TCF not only leads the reduction in the parameterisation error bound as delineated in Equation 23, but also lowers the distillation error bound.
@@ 4.3 Strategic Stochastic Sampling
It is worth noting that once our TCF is trained, one can use common PF-ODE solvers like DDIM or SDE-DPM-Solver to sample high quality images. But it is not easy to control the stochastic strength with existing solvers; hence, we proposed Strategic Stochastic Sampling for better stochasticity control.
Background.
Recalling the Multistep Consistency Sampling process in Algorithm 3, each iteration of Multistep Consistency Sampling from $t$ to $s$ can be expressed as:
Denoise: ${\bm{x}}_{t\rightarrow 0}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}({{\bm{x}}}_{t},t),$ (30)
Diffuse: $\hat{{\bm{x}}}_{s}\leftarrow\alpha_{s}{\bm{x}}_{t\rightarrow 0}+\sigma_{s}{\bm{z}},$
we found that the above process with parameterisation consistency function (Equation 14) can be rewritten as a DDIM sampling equation  (Song et al., 2020a) with $\eta=\sqrt{1-\alpha_{s}^{2}}$ in Equation 32:
${\bm{x}}_{s}=\alpha_{s}\underbrace{\left(\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\bm{\theta}}({\bm{x}}_{t})}{\alpha_{t}}\right)}_{\text{denoise}}+\underbrace{\sigma_{s}{\bm{z}}}_{\text{diffuse}}.$ (31)
Equation 31 leads us to review the general version of DDIM:
$\begin{split}{\bm{x}}_{s}&=\alpha_{s}{\left(\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\bm{\theta}}({\bm{x}}_{t})}{\alpha_{t}}\right)}\\ &\quad\quad\quad\quad+\sqrt{1-\alpha_{s}^{2}-\eta^{2}}\epsilon_{\bm{\theta}}({\bm{x}}_{t})+\eta{\bm{z}},\end{split}$ (32)
where occurs a parameter $\eta$ to control the stochasticity. And many related works (Karras et al., 2022; Song et al., 2020a; Zhang & Chen, 2022) observed that few stochasticity can help reduce the sampling error, but high stochasticity will lead to poor quality. Such stochasticity during multistep sampling will obviously degrade the final output.
Inspired by the ideas of stochasticity control term from DDIM and SDE-DPM-Solvers (Lu et al., 2022b), we propose Strategic Stochastic Sampling (SSS), which explicitly integrates the noise control term as an input parameter in the sampling process. Besides, our SSS could also be extended to different orders based on SDE-DPM-Solvers (Lu et al., 2022b). Here we demonstrate the 1-st order SSS solver, which is based on Equation 32:
$\begin{split}&{\bm{x}}_{s}=\alpha_{s}\left(\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\bm{\theta}}({\bm{x}}_{t})}{\alpha_{t}}\right)+\alpha_{s}\frac{\sigma_{s^{{}^{\prime}}}}{\alpha_{s^{{}^{\prime}}}}\epsilon_{\bm{\theta}}({\bm{x}}_{t})+\sqrt{1-\frac{\alpha_{s}^{2}}{\alpha_{s^{{}^{\prime}}}^{2}}}{\bm{z}}\\ &=\frac{\alpha_{s}}{\alpha_{s^{{}^{\prime}}}}\underbrace{\left(\alpha_{s^{{}^{\prime}}}\frac{{\bm{x}}_{t}-\sigma_{t}\epsilon_{\bm{\theta}}({\bm{x}}_{t})}{\alpha_{t}}+\sigma_{s^{{}^{\prime}}}\epsilon_{\bm{\theta}}({\bm{x}}_{t})\right)}_{\text{predicted}\;{\bm{x}}_{s^{{}^{\prime}}}={\bm{f}}({\bm{x}}_{t},t,s^{{}^{\prime}})}+\underbrace{\sqrt{1-\frac{\alpha_{s}^{2}}{\alpha_{s^{{}^{\prime}}}^{2}}}{\bm{z}}}_{\text{controllable noise}},\end{split}$ (33)
instead of $\eta$ in Equation 32, we introduced an stochastic parameter $\gamma$ that allows the control of the stochasticity in every step. As previously noted, the SSS framework is derived from the parameterisation of TCF and possesses the capability to explicitly manage stochastic intensities. Consequently, we define $\gamma$ as the parameterisation of target timestep $s$, with the linear range between full stochasticity and determinacy, thus $\gamma=1-s^{\prime}/s$ and $\eta=\sqrt{1-{\alpha_{s}^{2}}/{\alpha_{s^{{}^{\prime}}}^{2}}}$.
Sampling Process.
Similar to Multistep consistency sampling, we can simply decompose Equation 33 into the format of multi-step sampling, which is demonstrated in Algorithm 4:
Denoise: ${\bm{x}}_{t\rightarrow s^{\prime}}\leftarrow{\bm{f}}_{\bm{\theta^{*}}}^{\rightarrow s^{\prime}}({\bm{x}}_{t},t)$ (34)
Diffuse: $\hat{{\bm{x}}}_{s}\leftarrow\frac{\alpha_{s}}{\alpha_{s^{\prime}}}{\bm{x}}_{t\rightarrow s^{\prime}}+\sqrt{1-\frac{\alpha_{s}^{2}}{\alpha_{s^{\prime}}^{2}}}{\bm{z}}$ (35)
@@
Specifically, every sampling step in SSS includes the denoise sub-step according to the ODE solver and the diffuse sub-step based on Langevin SDE. In comparison with multistep consistency sampling, where the endpoint and noise level are fixed, SSS introduces the additional parameter $\gamma$ to control the destination point for the denoise step and allows for the adjustment of random noise level for the diffuse step, as detailed in Figure 2(b) and Algorithm 4.
In (Kim et al., 2023), the authors proposed a unified sampling method and have delicately crafted a $\gamma$-sampler for its $G_{\theta}$ function to leverage its ability for “anytime to anytime” prediction. This sampler incorporates a $\gamma$ parameter to regulate the decoder $G_{\theta}$ target, yet further endows it with the properties of both EDM sampler and deterministic sampler within the Multistep Consistency Sampling protocol. Another key finding in (Kim et al., 2023) is the analysis of the error bound in $\gamma$-sampling $\mathcal{O}(\sum_{n=0}^{N-1}\sqrt{t_{n}-\sqrt{1-\gamma^{2}}t_{n+1}})$, which aligns with the previous practical observation (Karras et al., 2022): less stochastic will lead to a better sampling quality.
It is worth noting that SSS was specially designed to control the stochasticity in LCM sampling with the well-defined $\gamma$ drived from DDIM sampling. Additionally, its alternating sampling framework is grounded on multistep consistency sampling, similar to the $\gamma$ sampling in (Kim et al., 2023). From this perspective, we notice that SSS is also capable of changing the stochastic intensity during its sampling procedure and has the same error bound as stated in (Kim et al., 2023). In another word, our method is also able to reduce the errors that arise from the $n$-th step to $\mathcal{O}(\sqrt{\tau_{n}-(1-\gamma)\tau_{(n+1)}})$ when $n\in[\![1,N-1]\!]$ and ultimately optimize the accumulated error. It is worth noting that in the process of sampling with TCF(1) and TCF(2), there is another discretization error $\mathcal{O}(h_{\Delta\tau}^{k+1})$ that cannot be eliminated, which is brought by the parameterisation of TCF as discussed in Equation 23.
We would like to highlight that the parameterisation form of TCF makes it versatile for community samplers, and our SSS provides an additional method to explicitly control the stochastic level for LCM. It is also worth noting that when $\gamma$ is low, the estimation error plays a more important role (Lyu et al., 2023). Hence, the optimal value of $\gamma$ should be determined empirically, as we show in Section 5.3.
@@
Figure 4: Qualitative comparison. For each prompt, images are generated using the same random seed for every model, without any cherry-picking. More results are provided in Section D.3.
@@
Figure 5: Qualitative effects of stochastic parameter $\gamma$ with same NFEs. Images under the same prompt but with different $\gamma$ applied during sampling. The leftmost image is sampled from LCM (Luo et al., 2023b). More samples can be found in Section D.2.
@@
4.4 Extension to Large Text Conditional Models
Trajectory consistency distillation can be considered as fine-tuning process, which directly occur on top of the pre-trained model. For scaling TCD to larger models (e.g., SDXL) with significantly reduced memory consumption, we incorporate Low-Rank Adaptation (LoRA) (Hu et al., 2021), a parameter-efficient fine-tuning method, into the distillation process. Additionally, the parameters of LoRA can be identified as a versatile acceleration module applicable to different fine-tuned models or LoRAs sharing the same base model without the need for additional training, aligning with the observations in (Luo et al., 2023b).
5 Experiments
5.1 Experimental Setup